There are 1000 spaces. Regular Parking is $5. Prefered parking is $10. They sold out $5750. How many of each type of parking are available?
$12 adult tickets were sold.
$7 student tickets were sold.
A total of 122 tickets were sold adding up to $1,099 total. How many student tickets were sold?

Answer by mananth(16946)

There are 1000 spaces. Regular Parking is $5. Prefered parking is $10. They sold out $5750. How many of each type of parking are available?
let regular parking spaces be x
prefered will be 1000-x
5x+10(1000-x)= 5750
5x+10000 - 10x= 5750
-5x=-4250
x= -4250/-5
x=850 regular spaces
balance will be 150 which is prefered space


$12 adult tickets were sold.
$7 student tickets were sold.
A total of 122 tickets were sold adding up to $1,099 total. How many student tickets were sold?
let student tickests sold be x
adult tickets sold will be 122-x
7x+12(122-x)=1099
7x+1464-12x=1099
-5x=-365
x=365/5
x= 73 tickets student tickets